HIGH-RESOLUTION NEAR-INFRARED SPECTROSCOPY OF H_{3}^{+}

\underline{CHRISTOPHER P. MORONG}, CHRISTOPHER F. NEESE and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

Last year we reported on the observation of 70 new rovibrational transitions of H_{3}^{+} in the near-infrared and visible regions.^{a} An additional 50 new lines have been observed since then including two hot band transitions in the nearinfrared region between 10,300-12,700 cm^{−1}. Both liquid nitrogen and water cooled discharges were used to study the ions. The rotational temperatures of the liquid nitrogen cooled discharge is about 700 K, while the water cooled discharge has a rotational temperature around 1000 K and was used for only a few of the strong high J >4 lines and the hot band transitions. These energy levels are above the barrier to linearity (>10,000 cm^{−1}), the regime in which H_{3}^{+} has enough energy to sample linear configurations. A high-resolution, high-sensitivity spectrometer based on a Ti:sapphire laser and incorporating velocity modulation and phase modulation with heterodyne detection^{b}, was used to observe the transitions. The transitions are more than 6200 times weaker than the fundamental band. Due to the abundance of strong hydrogen Rydberg transitions, both pure hydrogen and He/H_{2} plasmas were used to discriminate and identify the much weaker H_{3}^{+} transitions. The sparsity and weakness of the lines necessitated the use of the predicted intensities and frequencies^{c}^{d} to focus the wavelength region of our search. The measured rovibrational energy levels will assist in the development and verification of the theoretical calculations of H_{3}^{+} from first principles, which is the benchmark for ab initio theory.

aC. Morong, C. Neese, T. Oka, 61st Ohio State University International Symposium on Molecular Spectroscopy, TA03 (2006).

bJ. Gottfried, B. McCall, and T. Oka, J. Chem. Phys. 118, 10890 (2003).

cL. Neale, S. Miller, and J. Tennyson, Astrophys. J. 464, 516 (1996).

dP. Schiffels, A. Alijah, and J. Hinze, Mol. Phys. 101, 189 (2003).